Required information The following information applies to the questions displayed below.) Selected comparative financial stat2017 KORBIN COMPANY Comparative Balance Sheets December 31, 2019, 2018, and 2017 2019 2018 Assets Current assets $ 63,668 $ 4Required: 1. Complete the below table to calculate each years current ratio. Current Ratio Choose Denominator: Choose Numera2. Complete the below table to calculate income statement data in common-size percents. (Round your percentage answers to 2 d3. Complete the below table to calculate the balance sheet data in trend percents with 2017 as base year. (Round your percent
Answer #1
Answer 1
current ratio
numerator / denominator = Current Ratio
2019 63668 / 26296 =                   2.42 1
2018 42613 / 22293 =                   1.91 1
2017 56964 / 21692 =                   2.63 1
Answer 2
KORBIN COMPANY
Common-Size Comparative Statements
For Years ended December 31,2019,2018 and 2017
2019 % 2018 % 2017 %
Sales 100 % 100 % 100 %
Cost of goods sold 60.20% 62.70% 64.00%
Gross profit 39.80% 37.30% 36.00%
Selling expenses 14.20% 13.80% 13.20%
Administrative exps 9.00% 8.80% 8.30%
Total Expenses 23.20% 22.60% 21.50%
Income before taxes 16.60% 14.70% 14.50%
Income tax expenses 3.09% 3.01% 2.94%
Net Income 13.51% 11.69% 11.56%
Answer 3
KORBIN COMPANY
Balance Sheet data in Trend Percents
December 31,2019 ,2018 and 2017
2019 % 2018 % 2017 %
Assets
Current Assets 111.77% 74.81% 100.00%
Long term Investments 0.00% 29.48% 100.00%
Plant assets, net 185.05% 168.15% 100.00%
Total Assets 145.30% 120.70% 100.00%
Liabilities and equity
Current liabilities 121.22% 102.77% 100.00%
Common stock 140.00% 140.00% 100.00%
other paid in capital 157.50% 157.50% 100.00%
Retained earnings 158.69% 108.00% 100.00%
Total liabilities and capital 145.30% 120.70% 100.00%